A triode passes a plate current of 5mA at plate voltage 150V and grid voltage of -2V. If the grid voltage is changed to -35V, the plate current drops to 3.2mA but can be restored to 5mA by increasing plate voltage to 195V. Calculate

(1) mutual conductance

(2) a.c. plate resistance

(3) amplification factor


1
Expert's answer
2022-02-15T09:58:30-0500

1)

"g=\frac{I_{out}}{U_{in}}=0.47~\Omega,"

2)

"i=\frac{U_1-U_2}{R_1-R_2}=4.7~mA,"

3)

"\gamma=\frac{U_{out}}{U_{in}+U^*}=1.22."